OpenStudy (anonymous):

A sail is in the form of a right triangle that is three times as high as it is wide. The sail is made from 24 square meters of material. What is its height?

OpenStudy (anonymous):

a sail is in the form of a right triangle that is three times as high as it is wide. the sail is made from 24 square meters of material. what is its height?

OpenStudy (anonymous):

 No the area is 1/2 ab = 1/2 (a)(3a) = 24 (given)

OpenStudy (anonymous):

3a^2 = 48 a^2 = 16 a=4 i hope its correct